OpenStudy (anonymous):

Solve using any method 2x^2 + x - 5 = 0

OpenStudy (anonymous):

2. Solve using any method. -x^2 + 2x - 2 = 0

OpenStudy (anonymous):

2x^2+x-5=0 x=(-1+/-sqrt(1+40)/4) x=(sqrt(41)-1)/4 or x=(-sqrt(41)-1)/4

OpenStudy (anonymous):

Oh okay! Thanks!